# reference.bib

### v1

@@ 8.3. Euler product expression for the Dirichlet series
The question of this section is to obtain an Euler product expression for the Dirichlet series of interest in each of the above cases and relate it to standard $L-$function attached to $F$. By Theorem 7.2, the first step is to determine $\zeta_{\xi,\mathcal{P}}(s)$. Hence, we need to compute the quantity:
@@
with $\xi=(1,\mathbf{0},1)^{t}$, as we have explained in Remark 7. The steps we follow are:
@@
Let us now deal with the specific examples we have. In the following, let
@@
    $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. Here, we have $PSQ=\textup{diag}(1,3)$, with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$. We then have
@@
    and after substituting, the congruence of interest becomes ($q=3$ here):
@@
    with $t_{1}\pmod{d}$ and $t_{2}\pmod{3d}$. Now, by the form of the equation, we get $3\mid t_{2}$, so the congruence becomes:
@@
    with $t_{1},t_{2}\pmod{d}$. Now, we have $n(\xi;d)=N(T;d)$, where the last quantity is defined as the number of solutions to the congruence $T[t]\equiv-1\pmod{d}$, with
@@
    But $N(T;d)$ is multiplicative in $d$. Let now $p\notin\mathcal{P}\cup\{2,3\}$, so that 22 22 has a multiplicative inverse $\pmod{p}$ and $p\nmid\det{T}$. Then, by [diagonalisation], we know that for each $k\geq 1$, there is a matrix $U_{k}$ such that $T[U_{k}]\equiv R\pmod{p^{k}}$ with some diagonal matrix $R$. By then setting $t\longmapsto U_{k}t$, we can still consider $t_{i}\pmod{p^{k}}$ for all $i$ (as the determinant of $U$ is non-zero $\pmod{p^{k}}$). We then have $N(T;p^{k})=N(R;p^{k})$, with $R$ diagonal. We can now count the number of solutions $N(R;p^{k})$ by [congruences, Proposition 4]. In particular, if $R=\textup{diag}(a_{1},a_{2})$, we have
    $N(R;p^{k})=p^{k}\left[1-\frac{1}{p}\left(\frac{-a_{1}a_{2}}{p}\right)\right],$
    because $c_{1}(-1)=1,c_{p}(-1)=-1,c_{p^{l}}(-1)=0$ for $l\geq 2$ in the notation there. But $\det{R}=a_{1}a_{2}$, and $\det{R}\equiv(\det{U_{k}})^{2}\det{T}\pmod{p}$, so
@@
    where $2u\equiv 1\pmod{p}$. Therefore, we obtain:
@@
    $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, we have $PSQ=\textup{diag}(1,15)$ with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$. By following the above steps, the congruence becomes:
@@
    with $t_{1},t_{2}\pmod{d}$ and with exact same reasoning as before, we get:
    $n(\xi;p^{k})=p^{k}\left[1-\frac{\chi_{S}(p)}{p}\right].$
    for all $p\notin\mathcal{P}\cup\{2,3,5\}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    By then substituting $s=P^{-1}t$ the congruence becomes:
@@
    with $t_{1},t_{2},t_{3}\pmod{d}$ and $t_{4}\pmod{5d}$. But, we see that $5\mid t_{4}$, so after setting $t_{4}\longmapsto 5t_{4}$, the congruence becomes:
@@
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right].$
    for all $p\notin\mathcal{P}\cup\{2,5\}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, using the same method, we obtain:
@@
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right].$
    for all $p\notin\mathcal{P}\cup\{2,3\}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    In the last two cases, we omit the calculations, but by applying the exact same reasoning as above, we will again get:
    $n(\xi;p^{k})=p^{k(n-1)}\left[1-\frac{\chi_{S}(p)}{p^{n/2}}\right].$
    for all $p$ not in $\mathcal{P}$, $p\neq 2$ and $p\nmid\det{S}$ and with $n(\xi;d)$ multiplicative in $d$. Here $n=6,8$.
In all of the above cases, we therefore obtain (here $\mathcal{P}$ is possibly enlargened but still finite):
$\zeta_{\xi,\mathcal{P}}(s)=\prod_{p\notin\mathcal{P}}\zeta_{\mathcal{P}}(s-(n-1))\zeta_{\mathcal{P}}(s-n/2+1,\chi_{S})^{-1}.$
We therefore arrive at the following Theorem:
@@
Let $S$ denote any of the matrices of Section 8.2 with even rank $n$ and $\mathcal{P}$ a sufficiently large finite set of primes, containing the primes described in Theorem 7.2 and the primes so that the conditions of [shimura_orthogonal, Proposition 5.13] are satisfied. Then, if we define
$\chi_{S}(p):=\left(\frac{(-1)^{n/2}\det{S}}{p}\right),$
we have
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)=A(\xi)\zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}\zeta_{\mathcal{P}}\left(s-k+\frac{n+4}{2},\chi_{S}\right)^{-1}L_{\mathcal{P}}\left(F;s-k+\frac{n+2}{2}\right)\times$
×$\times\prod_{i=1}^{n-1}\zeta_{\mathcal{P}}(s-k+(n+2)-i)^{-1},$
where again $\xi=(1,\mathbf{0},1)^{t}$.
@@
The proof follows from Theorem 7.2 after we choose $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. From [shimura_orthogonal, Proposition 5.15], we have that (since $\mathcal{P}$ is chosen so that conditions of [shimura_orthogonal, Proposition 5.13] are satisfied):
$L_{\mathcal{P}}\left(\mathbf{1},s-k+\frac{n+3}{2}\right)=\prod_{i=1}^{n}\zeta_{\mathcal{P}}(s-k+(n+2)-i),$
as in general $L_{\textup{Sug}}\left(\textbf{1};s-\frac{n-1}{2}\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L-$function we encounter in [sugano] and [shimura_orthogonal] respectively, as these are normalised differently. Finally, the computation of $\zeta_{\xi,\mathcal{P}}$ has been performed above. ∎
9. Conclusions
In this paper, our main goal has been to obtain Euler product expansions for the Dirichlet series $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$, extending in this way, the method of Kohnen and Skoruppa in [kohnen_skoruppa] to the orthogonal setting. However, it seems that our case presents some limitations, which do not allow us to give uniform results for orthogonal groups of arbitrary rank.
To be more specific, the first main difficulty arises from the fact that it does not seem to exist an obvious connection between $\sum_{i=1}^{h}\zeta_{\xi_{i}}(s-k+n+1)f_{j}(u_{i})$ and $L\left(\overline{f_{j}};s-k+(n+3)/2\right)$. This does not happen in Kohnen and Skoruppa, as this connection can be seen by using the correspondence between equivalence classes of binary quadratic forms with fixed discriminant and ideal classes in the class group of a quadratic extension. This has led us restricting our attention to the case where $h=1$.
In order therefore to find explicit examples, we need to pick $\xi$ such that the size of $H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C)$ is just 11 11. The choice of $\xi=(1,\textbf{0},1)^{t}$, apart from being natural and simple, satisfies another very important purpose. In particular, it gives us that any vector in $\Xi_{d}$ of Lemma 6.1 lies in the same equivalence class of $\xi$ under not only the action of $\Gamma(L_{0})$, but also of $\Gamma^{+}(L_{0})$. This is crucial for the explicit computation of $\zeta_{\xi,\mathcal{P}}(s)$, which is done in Section 8. In the rank 11 11 case, Shimura in [shimura_diophantine_2006] has given an explicit formula for the index $[H(\xi)_{\mathbb{A}}:H(\xi)_{\mathbb{Q}}(H(\xi)_{\mathbb{A}}\cap C)]$, which we use to find all the cases for which this index is 11 11 (see Section 8.1). In the rank $n\geq 2$, such a formula is not available, so we need an explicit description of $H(\xi)_{\mathbb{Q}}$ as a definite orthogonal group. It is well-known that there is a finite number of definite orthogonal groups with class number 11 11 and their rank is at most 1010 1010 (cf. [watson]). Using, therefore, the enumeration of definite orthogonal groups with class number 11 11 in [hanke2011enumerating] as well as [shimura2004arithmetic, Proposition 11.12], we arrive at the specific examples of Section 8 for rank $n\geq 2$.
[…]

### v2

@@ 9.3. Euler product expression for the Dirichlet series
The question of this Section is to obtain an Euler product expression for the Dirichlet series of interest in each of the above cases and relate it to standard $L$-function attached to $F$. Again, let $\mathcal{P}$ be as in Theorem 8.2, containing also the prime $2$. In particular, $\mathcal{P}$ contains the primes factors of $q$, hence of $\det S$ as well. By Theorem 8.2, the first step is to determine $\zeta_{\xi,\mathcal{P}}(s)$. Hence, we need to compute the quantity
@@
with $\xi=(1,\mathbf{0},1)^{t}$, as we have explained in the beginning of Section 9. The steps we follow are:
@@
Let us now deal with the specific examples we have. In the following, let for $p\not\in\mathcal{P}$
@@
    $S=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. We have $PSQ=\textup{diag}(1,3)$, with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$ and specified $Q$. We then have
@@
    and after substituting, the congruence of interest becomes ($q=3$ here)
@@
    with $t_{1}\pmod{d}$ and $t_{2}\pmod{3d}$. Now, by the form of the equation, we get $3\mid t_{2}$, so the congruence becomes
@@
    with $t_{1},t_{2}\pmod{d}$. Now, we have $n(\xi;d)=N(T;d)$, where the last quantity is defined as the number of solutions $t=(t_{1},t_{2})\pmod{d\mathbb{Z}^{2}}$ to the congruence $T[t]\equiv-1\pmod{d}$, with
@@
[…]
    $N(R;p^{k})=p^{k}\left[1-\frac{1}{p}\left(\frac{-a_{1}a_{2}}{p}\right)\right].$
    But $\det{R}=a_{1}a_{2}$, and $\det{R}\equiv(\det{U_{k}})^{2}\det{T}\pmod{p}$, so
@@
    where $2u\equiv 1\pmod{p}$. Therefore, we obtain
@@
    $S=\begin{pmatrix}2&-1\\ -1&8\end{pmatrix}$. In this case, we have $PSQ=\textup{diag}(1,15)$ with $P=\begin{pmatrix}1&1\\ 1&2\end{pmatrix}$ and specified $Q$. By following the above steps, the congruence becomes
@@
    with $t_{1},t_{2}\pmod{d}$. With exact same reasoning as before, we get
    $n(\xi;p^{k})=p^{k}\left[1-\frac{\chi_{S}(p)}{p}\right],$
    for all $p\notin\mathcal{P}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    By then substituting $s=P^{-1}t$ the congruence becomes
@@
    with $t_{1},t_{2},t_{3}\pmod{d}$ and $t_{4}\pmod{5d}$. But, again, $5\mid t_{4}$, so after setting $t_{4}\longmapsto 5t_{4}$, we have
@@
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right],$
    for all $p\notin\mathcal{P}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    $S=\begin{pmatrix}2&-1&0&0\\ -1&2&0&0\\ 0&0&2&-1\\ 0&0&-1&2\end{pmatrix}$. In this case, using the same method, we obtain
@@
    $n(\xi;p^{k})=p^{3k}\left[1-\frac{\chi_{S}(p)}{p^{2}}\right],$
    for all $p\notin\mathcal{P}$ with $n(\xi;d)$ multiplicative in $d$.
@@
    In the last two cases, we omit the calculations, but by applying the exact same reasoning as above, we will again get
    $n(\xi;p^{k})=p^{k(n-1)}\left[1-\frac{\chi_{S}(p)}{p^{n/2}}\right],$
    for all $p\not\in\mathcal{P}$ and with $n(\xi;d)$ multiplicative in $d$. Here $n=6,8$.
In all of the above cases, we therefore obtain (here $\mathcal{P}$ can be possibly enlarged but still finite)
$\zeta_{\xi,\mathcal{P}}(s)=\prod_{p\notin\mathcal{P}}\zeta_{\mathcal{P}}(s-(n-1))\zeta_{\mathcal{P}}(s-n/2+1,\chi_{S})^{-1},$ (22)
where $\zeta_{\mathcal{P}}(s,\chi_{S}):=\prod_{p\not\in\mathcal{P}}(1-\chi_{S}(p)p^{-s})^{-1}$. We therefore arrive at the following Theorem.
@@
Let $S$ denote any of the matrices of Section 9.2 with even rank $n$ and $\mathcal{P}$ a finite set of primes, containing the primes described in the proof of Theorem 8.2, the prime $2$, and the primes so that the conditions of [23, Proposition 5.13] are satisfied for all $p\not\in\mathcal{P}$. For all $p\not\in\mathcal{P}$, we define
$\chi_{S}(p):=\left(\frac{(-1)^{n/2}\det{S}}{p}\right).$
If $F$ and $\mathcal{P}_{k,-q,r}$ are as in Theorem 8.2 and $\xi=(1,\mathbf{0},1)^{t}$ (in particular $A(\xi)\neq 0$), we have
$\mathcal{D}_{F,\mathcal{P}_{k,-q,r},\mathcal{P}}(s)=A(\xi)\zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}\zeta_{\mathcal{P}}\left(s-k+(n+4)/2,\chi_{S}\right)^{-1}\times\\ \times L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\prod_{i=1}^{n-1}\zeta_{\mathcal{P}}(s-k+(n+2)-i)^{-1}.$
@@
The proof follows from Theorem 8.2 after we choose $f$ to be the constant 1 on $H(\xi)_{\mathbb{A}}$. From [23, Proposition 5.15], we have that (since $\mathcal{P}$ is chosen so that conditions of [23, Proposition 5.13] are satisfied):
$L_{\mathcal{P}}\left(\mathbf{1},s-k+(n+3)/2\right)=\prod_{i=1}^{n}\zeta_{\mathcal{P}}(s-k+(n+2)-i),$
as in general $L_{\textup{Sug}}\left(\textbf{1};s-(n-1)/2\right)=L_{\textup{Shi}}(s)$. Here, the notation means the $L$-function we encounter in [27] and [23] respectively, as these are normalised differently. Finally, the computation of $\zeta_{\xi,\mathcal{P}}(s)$ has been performed in (22). ∎
Funding: This work was supported by the HIMR/UKRI "Additional Funding Programme for Mathematical Sciences", grant number EP/V521917/1, as well as by a scholarship from the Onassis Foundation.